OpenStudy (anonymous):

Sam has a total of 58 DVD's and CD's. If the number of CD's is two more than three times the number of DVD's, how many CD's does he have?

OpenStudy (radar):

Let x = number of CDs 58-x = number of DVDs 3(58-x)+2 = three times the DVDs +2 or 3(58-x)+2 =x 174-3x +2 = x 176=4x x=44 Sam has 44 CDs and 14 DVD's